OpenStudy (anonymous):

a particle starts SHM from the mean position. its amplitude is A and time is T.at what time when its speed is half of the maximum speed, its displacement is y is?

OpenStudy (jamesj):

You mean its period is T?

OpenStudy (anonymous):

here i did v^2=A^2-X^2 and here we know max velocity is Aw=vmax w=2pie/T but nw i m unable to solve it further by keeping its value in the above formula

OpenStudy (anonymous):

A2pie/T=A^2-X^2 its complex ... am i doing sm mistake because its becoming complex means i had done some mistakes anywhere

OpenStudy (jamesj):

Please clarify. T is the period of the system?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

hey for pie^2 we can take it approx 10 rit

OpenStudy (anonymous):

as its value is 3.14 so by squaring we can take approx 10??

OpenStudy (jamesj):

Ok then. In which case the angular frequency \( \omega \) is given by \[ \omega = \frac{2\pi}{T} \] and the equation of motion of the system is \[ x(t) = A \cos( \omega t + \phi) \] Now, to solve the problem, first find \( \phi \) and then use this equation to find the value of \( t \) for which \( v(t) = \dot{x}(t) \) is half the maximum velocity, which you'll also want to find.

OpenStudy (anonymous):

how can we find phi will u plz solve dis ques i think i need a solution here to see where i m lacking

OpenStudy (jamesj):

Look at the wording of the question and ask yourself what happens when t = 0.

OpenStudy (anonymous):

wait a min they are asking time t rit or displacement??

OpenStudy (jamesj):

Both I think. This is where writing the question in clear English is so important.

OpenStudy (anonymous):

in my text they always write in a twisted form and i cant fight with education system alone to write it in good english :P

OpenStudy (anonymous):

hey i just rechecked the formula i used it was wrong the formula must be v^2=w^2A^2-A^2x^2 my mistake i got my ans now :P

OpenStudy (jamesj):

To find phi: we're told that "a particle starts SHM from the mean position" hence x(t=0) = 0. But \[ x(0) = A \cos(0 + \phi) \] thus \( \cos \phi = 0 \) and therefore \( \phi = \pi/2 \). We can now write \[ x(t) = A\cos(\omega t + \phi) = A\sin(\omega t) \]

OpenStudy (anonymous):

i got it i got it nw dont solve plz......

OpenStudy (jamesj):

I'm solving from first principles. But I'll stop if you really want.

OpenStudy (anonymous):

yea plz actually if i see solution i dont use my mind dat is bad u cant help me in exam time dats y ...

OpenStudy (jamesj):

I respect that.

OpenStudy (anonymous):

thanQ listen i m in confusion wid other i m nt able to make eaquation plz help me